# EE-MLLM: A Data-Efficient and Compute-Efficient Multimodal Large Language Model

### v1

@@ Abstract
In the realm of multimodal research, numerous studies leverage substantial image-text pairs to conduct modal alignment learning, transforming Large Language Models (LLMs) into Multimodal LLMs and excelling in a variety of visual-language tasks. The prevailing methodologies primarily fall into two categories: self-attention-based and cross-attention-based methods. While self-attention-based methods offer superior data efficiency due to their simple MLP architecture, they often suffer from lower computational efficiency due to concatenating visual and textual tokens as input for LLM. Conversely, cross-attention-based methods, although less data-efficient due to additional learnable parameters, exhibit higher computational efficiency by avoiding long sequence input for LLM. To address these trade-offs, we introduce the Data-Efficient and Compute-Efficient Multimodal Large Language Model (EE-MLLM). Without introducing additional modules or learnable parameters, EE-MLLM achieves both data and compute efficiency. Specifically, we modify the original self-attention mechanism in MLLM to a composite attention mechanism. This mechanism has two key characteristics: 1) Eliminating the computational overhead of self-attention within visual tokens to achieve compute efficiency, and 2) Reusing the weights on each layer of LLM to facilitate effective modality alignment between vision and language for data efficiency. Experimental results demonstrate the effectiveness of EE-MLLM across a range of benchmarks, including general-purpose datasets like MMBench and SeedBench, as well as fine-grained tasks such as TextVQA and DocVQA.
@@
Introduction
Recently, Large Language Models (LLMs) have gained significant attention due to their surprising performance on various natural language tasks. However, real-world scenarios often involve more than just linguistic modality, making it crucial to extend LLM to Multimodal LLMs. The key to expansion is to perform modality alignment, that is, to learn to map the remaining modalities with the same semantics to their corresponding language modality in the feature space of the pre-trained LLM.
In the field of multimodal LLM research involving vision and language, a series of research works on modality alignment have been carried out and achieved excellent performance. These methods can be roughly divided into two categories: self-attention-based methods and cross-attention-based methods.
The self-attention-based methods firstly leverage a simple MLP layer to project the visual tokens into the input space of LLM, and then concatenate the visual tokens with the textual tokens, which is shown in Fig.1 (1). These flattened concatenated long token sequences are fed into the LLM to further model the dependencies of various tokens. These methods exhibit great data efficiency due to two factors: 1) The alignment module is simple and has a small number of learnable parameters. 2) Visual tokens aligned in the input space of LLM will naturally lead to the alignment between vision and language in each layer of LLM. Therefore, the training procedure does not need to focus on the modality alignment within the LLM, enabling data efficiency. For example, LLaVA (Liu et al. 2023b) achieves promising performance by pre-training on only 558k data and fine-tuning on only 665k instruction data. However, the flattening operation of directly concatenating visual tokens and textual tokens expands the length of the token sequence from $T$ to $V+T$ (where $T$ is the number of textual tokens, $V$ is the number of visual tokens), and the computational cost of the language model layer (self-attention structure) increases from $O(T^{2})$ to $O((V+T)^{2})$, results in lower computer-efficiency. Especially when the input image is high-resolution, the number of visual tokens increases, and the compute inefficiency will prolong the training and inference time. Specifically, LLaVA (Liu et al. 2023b) and Deepseek-VL (Lu et al. 2024) have 576 visual tokens, Sphinx (Lin et al. 2023) utilizes 2,890 visual tokens.
@@ Introduction
The cross-attention-based methods directly insert new cross-attention layers between existing pre-trained language model layers. As shown in Fig.1 (2), the keys and values in these layers are obtained from the visual tokens while the queries are derived from the text tokens. These methods have higher compute efficiency because the length of the input sequence is always kept as $T$ and does not increase with the increase of visual tokens. However, these methods usually require a large amount of pre-training data, leading to data inefficiency. For example, Flamingo (Alayrac et al. 2022) is pre-trained on 1.8 billion image-text pairs (Jia et al. 2021) and interleaved data. EVLM (Chen et al. 2024) is pre-trained on 2.5 billion image-text pairs and 50 million interleaved image-text data. This data inefficiency arises from two factors: 1) The additional cross-attention layers will introduce a large number of learnable parameters, necessitating more training data. 2) The training procedure has to focus on the modality alignment between vision and language on different layers of LLM, posing a huge optimization complexity for modality alignment.
@@
By reviewing existing methods, we conclude that the key points to ensure data efficiency and compute efficiency are 1) Avoiding an increase in the length of input sequence for LLM, 2) The modality alignment module should have fewer parameters, and the alignment needs to be effective. Taking two summary viewpoints into account, we present a data-Efficient and compute-Efficient Multimodal Large Language Model (EE-MLLM). Specifically, we propose a composite attention mechanism to replace the original self-attention mechanism in MLLM. The composite attention mechanism has two features: 1) Eliminating the computational overhead of self-attention within visual tokens to achieve compute efficiency. The original self-attention mechanism of MLLM consists of three components: self-attention within visual tokens, self-attention within text tokens, and cross-attention between text and visual tokens. We observe that the self-attention within visual tokens is redundant. This is because a) the interaction between visual tokens is already well-processed by the vision encoder, and b) LLM can facilitate the implicit interaction among visual tokens via the information aggregation property of LLM (Wang et al. 2023b). Therefore, we eliminate the self-attention within visual tokens and maintain the length of the input sequence for LLM as the length of text tokens, thereby facilitating computational efficiency. 2) Instead of introducing additional modules like cross-attention-based methods, we reuse the weights on each layer of LLM to facilitate effective modality alignment between vision and language for data efficiency. As a result, without introducing additional parameters, visual tokens aligned in the input space of LLM can naturally facilitate the alignment between vision and language in each layer of LLM. This means the training procedure does not need to focus on the modality alignment within the LLM, which enables data efficiency.
To demonstrate the effectiveness of EE-MLLM, we evaluate its performance on both general benchmarks and fine-grained benchmarks, such as MMBench-EN, MME, ScienceQA, AI2D, TextVQA, and DocVQA. In comparison with state-of-the-art MLLMs, our EE-MLLM exhibits promising results on both general and fine-grained benchmarks. Simultaneously, EE-MLLM proves to be computationally efficient in the inference stage. When applied to high-resolution image inputs, EE-MLLM maintains comparable performance at a significantly reduced computational cost. For instance, with a $980\times 980$ input image, the number of visual tokens can reach 4,900, and the FLOPs of EE-MLLM are $70\%$ of those in self-attention-based methods. We also assess the inference speed of EE-MLLM in a real-world scenario, confirming its efficiency. Specifically, when the number of generated tokens is set to 32, EE-MLLM achieves an inference speed of 77 tokens per second on a single NVIDIA H800 GPU, which is 1.9 times faster than self-attention-based methods at the same resolution $980\times 980$.
@@
    We revisit modality alignment in multimodal LLMs, pointing out the issue of efficiency and efficacy in self-attention-based and cross-attention-based methods.
@@
    We propose a composite attention mechanism for EE-MLLM to enhance the data and computational efficiency.
@@
    Our EE-MLLM achieves outstanding performance on a variety of benchmarks, while the inference speed is much improved.
Related Work
@@
Self-attention-based methods. Self-attention-based methods (Dai et al. 2024; Zhu et al. 2024; Chen et al. 2023a; Bai et al. 2023; Chen et al. 2023b; Fang et al. 2023) employ a projector to align visual features with the input space of LLM. LLaVA (Liu et al. 2023b, a) uses a simple MLP layer to connect the vision encoder of CLIP (Radford et al. 2021) with Vicuna (Zheng et al. 2023). DeepSeek-VL (Lu et al. 2024) employs hybrid vision encoders, including SigLIP (Zhai et al. 2023) for high-level semantic features and SAM-B (Kirillov et al. 2023) for low-level features. Considering the scenarios that require fine-grained information, such as document recognition and form understanding, recent work extends MLLM to support high-resolution inputs. LLaVA-Next (Liu et al. 2024) proposes dividing the image into small patches for extracting local features and concatenating the downsampled image feature as global context. Monkey (Li et al. 2024b) also partitions high-resolution images into small patches, processing each patch independently with a trainable resampler. Since the self-attention-based methods are compute-inefficiency, the overhead for high-resolution inputs can be more severe. Deco (Yao et al. 2024) and HoneyBee (Cha et al. 2024) carefully design projectors to reduce the number of visual tokens with minimal information loss. However, these methods still compromise fine-grained capabilities when reducing the number of visual tokens. Our EE-MLLM, a cross-attention-based architecture, achieves state-of-the-art performance while maintaining computational efficiency.
@@
Cross-attention-based methods. Cross-attention-based methods (Laurençon et al. 2023; Alayrac et al. 2022; Awadalla et al. 2023; Chen et al. 2024) typically incorporate additional cross-attention modules into the decoder layers of LLM to integrate visual information without extending the length of the text token sequence. Flamingo (Alayrac et al. 2022) introduces a novel gated xattn-dense block, incorporating a tanh-gating mechanism for training stability. Flamingo is trained on large-scale interleaved image-text data and supports in-context few-shot learning. EVLM (Chen et al. 2024) enhances Flamingo by: (1) utilizing hierarchical ViT features without a resampler, (2) replacing the image token with a set of learnable tokens in the text input, and (3) introducing MoE to boost performance. Although these cross-attention-based methods are computationally efficient, the additional cross-attention blocks lead to a large number of trainable parameters and a heavy reliance on pre-training data. Our EE-MLLM avoids introducing any extra trainable parameters and repurposes the weights of LLM to map visual tokens into the LLM feature space, thus achieving promising performance with a small amount of training data.
Method
Model Architecture
@@
EE-MLLM comprises a visual encoder, a two-layer MLP serving as the projector, and the LLM with composite decoder layers. We propose a composite attention mechanism and design a composite decoder layer for EE-MLLM to achieve data efficiency and compute efficiency. The composite decoder layer consists of an aligner and a composite attention module.
@@
Composite Attention Module. The original self-attention mechanism comprises self-attention within visual tokens, self-attention within text tokens, and cross-attention between text tokens and visual tokens. We observe that self-attention within visual tokens is redundant for two reasons: 1) The interaction between visual tokens is already well-learned in the vision encoder, and 2) LLM can enable the implicit interaction between visual tokens via the information aggregation property (Wang et al. 2023b). Therefore, we eliminate self-attention within visual tokens and obtain the composite attention module. We assume that visual tokens $I\in\mathbb{R}^{k\times h}$ and text tokens $T\in\mathbb{R}^{n\times h}$, where $k$ and $n$ are the lengths of visual tokens and text tokens, respectively. And $h$ denotes the dimension of hidden states. The original self-attention can be formulated as follows:

### v3

@@ Abstract
Recent advancements in Multimodal Large Language Models (MLLMs) have demonstrated satisfactory performance across various vision-language tasks. Current approaches for vision and language interaction fall into two categories: self-attention-based and cross-attention-based methods. However, both approaches present inherent limitations, forcing a trade-off between data and computational efficiency. To address this issue, we introduce the Data-Efficient and Compute-Efficient MLLM (EE-MLLM). Specifically, we modify the original self-attention mechanism in MLLM to a composite attention mechanism. This mechanism has two key characteristics: 1) eliminating the computational overhead of self-attention among visual tokens to achieve compute efficiency, and 2) reusing the weights from each layer of LLM to facilitate effective vision-language modality alignment for data efficiency. As a result, EE-MLLM significantly outperforms Flamingo with limited training data, and reduces the prefilling time to 79 ms on an H800 GPU, compared to LLaVA’s 277 ms. To further investigate the efficiency of EE-MLLM, we present a training-free variant named EE-MLLM-F, which reduces the computation cost of self-attention-based method without additional training. Experimental results demonstrate the effectiveness of EE-MLLM across a range of benchmarks, including general-purpose datasets like MMBench and SeedBench, as well as fine-grained tasks such as TextVQA and DocVQA.
@@
1 Introduction
Multimodal Large Language Models (MLLMs) have gained significant attention due to their remarkable performance on diverse vision-language tasks, including image captioning, Visual Question Answering (VQA) and so on. Currently, mainstream MLLMs can be categorized into self-attention-based and cross-attention-based methods. However, these two predominant approaches entail a trade-off between compute and data efficiency. In general, cross-attention-based methods excel in computational efficiency but require substantial training data for effective modality alignment, thus exhibiting low data efficiency. Conversely, self-attention-based methods have high data efficiency, which can achieve modality alignment with limited data, but often suffer from high computational complexity, resulting in reduced computational efficiency.
@@ 1 Introduction
Specifically, self-attention-based methods project visual tokens into LLM’s input space with a simple MLP and concatenate them with text tokens, as illustrated in Fig.1 (1). Their data efficiency stems from: 1) Few learnable parameters in the alignment module, and 2) When visual tokens are aligned to the input space via a projector during input, these remain consistent with the processing space of each LLM’s decoder layer. Therefore, LLaVA achieves strong performance with only 558k pre-training and 665k instruction-tuning samples [28], in other words, it does not require a large amount of data for modality alignment. Cross-attention-based methods insert cross-attention layers between existing LLM layers. As shown in Fig.1 (2), the keys and values in these layers are derived from visual tokens, while the queries come from text tokens. These methods introduce a large number of learnable parameters, which require substantial training data, leading to data inefficiency. For example, Flamingo is pre-trained on 1.8 billion image-text pairs [1, 16] and interleaved data. Thus, this data inefficiency arises from two factors: 1) Learnable parameters from additional cross-attention layers, and 2) The training procedure has to focus on the modality alignment between vision and language on different layers of LLM, posing a huge optimization complexity for modality alignment.
Regarding compute efficiency, self-attention methods concatenate visual and text tokens, which increases sequence length from $T$ to $V+T$, raising computational complexity of LLM from $O(T^{2})$ to $O((V+T)^{2})$. In contrast, cross-attention methods can achieve high compute efficiency because the input sequence length remains fixed at $T$ and does not increase with the number of visual tokens.
@@
By reviewing existing methods, we conclude that the key points to ensure compute efficiency and data efficiency are 1) Avoiding an increase in the length of input sequence for LLM, 2) The modality alignment module should have fewer parameters, and the alignment needs to be effective. Taking two summary viewpoints into account, we present a data-Efficient and compute-Efficient Multimodal Large Language Model (EE-MLLM). Specifically, we propose a composite attention mechanism to replace the original self-attention mechanism in LLM. The composite attention mechanism has two features: 1) Eliminating the computational overhead of self-attention within visual tokens to achieve compute efficiency. Similar to FastV [9], we observe that the self-attention within visual tokens is redundant. This redundancy stems from: a) the interaction between visual tokens is already well-processed by the vision encoder, and b) LLM can facilitate the implicit interaction among visual tokens via the information aggregation property of LLM [42]. Therefore, we eliminate self-attention among visual tokens and maintain the length of the input sequence for the LLM to be the same as that of text tokens, thereby facilitating computational efficiency. 2) Instead of introducing additional modules like cross-attention-based methods, we reuse the weights on each layer of LLM to facilitate effective modality alignment between vision and language for data efficiency. As a result, without introducing additional parameters, visual tokens aligned in the input space of LLM can naturally facilitate the alignment between vision and language in each layer of LLM. This means the training procedure does not need to focus on the modality alignment within the LLM, which enables data efficiency. With the proposed composite attention, EE-MLLM can achieve data-efficient and compute-efficient at the same time. We also investigate a training-free variant of EE-MLLM where we directly apply the composite attention mechanism to select layers of LLaVA, thereby reducing computational cost without requiring additional training. Notably, this approach maintains comparable performance, demonstrating computational efficiency without sacrificing accuracy.
We evaluate EE-MLLM on both general benchmarks and fine-grained benchmarks and our model demonstrates promising results across all these evaluations while maintaining high computational efficiency during inference. For a $980\times 980$ input image, EE-MLLM requires only $76\%$ of the FLOPs needed by self-attention methods. Moreover, in real-time scenarios that require high response speed but fewer output tokens, EE-MLLM can significantly reduce the prefilling time. On a single NVIDIA H800 GPU, EE-MLLM takes only 79 ms for prefilling with one input image, whereas LLaVA consumes 277 ms.
@@
    We propose composite attention mechanism for EE-MLLM, which enhances both data and compute efficiency by eliminating redundant attention calculation among visual tokens and reusing existing LLM weights for modality alignment.
@@
    We also introduce a training-free variant EE-MLLM-F, which adapts the composite attention mechanism to existing self-attention-based models, reducing computational costs without requiring additional training while preserving performance.
@@
    Experimental results demonstrate that EE-MLLM achieves superior performance across various benchmarks, while requiring only 76% of the FLOPs compared to self-attention methods and significantly reducing the prefilling time from 277 ms to 79 ms.
2 Related Work
@@
Self-attention-based methods. Self-attention-based methods [10, 48, 6, 3, 8, 12, 33] employ a projector to align visual features with the input space of LLM. LLaVA [28, 27] uses a simple MLP layer to connect the vision encoder of CLIP [39] with Vicuna [47]. Considering the scenarios that require fine-grained information, such as document recognition and form understanding, recent work extends MLLM to support high-resolution inputs, resulting in long input sequences with numerous visual tokens. LLaVA-Next [29] proposes dividing the image into small patches for extracting local features and concatenating the downsampled image feature as global context. Monkey [26] also partitions high-resolution images into small patches, processing each patch independently with a trainable resampler. Another line of research focuses on compressing the number of visual tokens to reduce computational cost. Deco [44] and HoneyBee [5] carefully design projectors to reduce the number of visual tokens with minimal information loss. However, these methods still compromise fine-grained capabilities when reducing the number of visual tokens. Our EE-MLLM, a cross-attention-based architecture, achieves state-of-the-art performance while maintaining computational efficiency.
@@
Cross-attention-based methods. Cross-attention-based methods [22, 1, 2, 7] typically incorporate additional cross-attention modules into the decoder layers of LLM to integrate visual information without extending the length of the text token sequence. Flamingo [1] introduces a novel gated xattn-dense block, incorporating a tanh-gating mechanism for training stability. Flamingo is trained on large-scale interleaved image-text data and supports in-context few-shot learning. EVLM [7] enhances Flamingo by: (1) utilizing hierarchical ViT features without a resampler, (2) replacing the image token with a set of learnable tokens in the text input, and (3) introducing MoE to boost performance. Although these cross-attention-based methods are computationally efficient, the additional cross-attention blocks lead to a large number of trainable parameters and a heavy reliance on pre-training data. Our EE-MLLM avoids introducing any extra trainable parameters and repurposes the weights of LLM to map visual tokens into the LLM feature space, thus achieving promising performance with a small amount of training data.
3 Method
3.1 Model Architecture
@@
EE-MLLM is composed of three main components: a visual encoder, a projector made up of two-layer MLPs, and an LLM with composite decoder layers. The composite decoder layer contains an aligner and a composite attention module, designed to ensure data and compute efficiency.
3.1.1 Composite Attention Module
The original self-attention mechanism builds correlations in three aspects: intra-interaction among visual tokens, intra-interaction among text tokens, and inter-interaction among visual tokens and text tokens. In decoder layers, we observe that the process with visual tokens is redundant for two reasons: 1) The interaction among visual tokens is already well-learned in the vision encoder, and 2) the visual information has been aggregated into the text tokens through the inter-interaction among text tokens and visual tokens [42]. Therefore, we eliminate intra-interaction among visual tokens and obtain the composite attention module.
@@ 3.1.1 Composite Attention Module
We assume that visual tokens $I\in\mathbb{R}^{k\times h}$ and text tokens $T\in\mathbb{R}^{n\times h}$, where $k$ and $n$ are the lengths of visual tokens and text tokens, respectively. And $h$ denotes the dimension of hidden states. The original self-attention can be formulated as follows:
@@
where $X^{q}=[I;T]W_{q}$, $X^{k}=[I;T]W_{k}$, $X^{v}=[I;T]W_{v}$, $W_{q}$, $W_{k}$ and $W_{v}$ are the weights of the self-attention module. The symbol [;][;][ ; ] denotes the concatenation operation of two variables in the sequence dimension.